OpenStudy (anonymous):

May someone please help me? Assuming the x > 0 which of these expressions is equivilent to 11 sqrt 245x^3 + 9 sqrt 45x^3? a. 5x sqrt 104x b. 20 sqrt 290x^6 c. 20x sqrt 290x d. 104x sqrt 5x

OpenStudy (shayaan_mustafa):

Hi the-soulseeker-vio :) How are you?

OpenStudy (shayaan_mustafa):

I am having difficulty in reading your expression kindly write them by using equation editor dear.

OpenStudy (anonymous):

\[11\sqrt{245^3} + 9\sqrt{45x^3}\]

OpenStudy (anonymous):

I had a typo: \[11\sqrt{245x^{3}}+9\sqrt{45x^{3}}\]

OpenStudy (anonymous):

Yes. That is it.

OpenStudy (anonymous):

Step 1: Factor out what's under each square root. Remember: \[\sqrt{a^{2}b} =a \sqrt{b}\]

OpenStudy (anonymous):

Ok, got that.

OpenStudy (radar):

Yes, I ooopsed, should of been a 5 * 49,,,, I deleted it. Let me try again with more steps, just the first term only \[11\sqrt{245x ^{3}}=11\sqrt{5*49 x ^{2}x}=77x \sqrt{5x}\] did@the-soulseeker-vio you get that?

OpenStudy (radar):

Thanks, @Wired,

OpenStudy (radar):

Looks like I got a monolog going lol.

OpenStudy (anonymous):

He closed the questionand already posted 2 others, so I'm assuming he solved it.

OpenStudy (radar):

O.K. good. A medal for your guidance for both the-soulseeker=vio and radar lol.